I want to know how to solve the following system using process of elimination please:
2y + 4z = 2
x + y + 2z = -2
2x + y + z = 2

Steps on how to solve this would be wonderful!!

Answer by mananth(16949)

2y + 4z = 2
2y=2-4z
y=1-2z

x + y + 2z = -2
substitute y in the equation
x+1-2z+2z=-2
x+1=-2
x=-3

2x + y + z = 2
substitute x & z in the equation
-6+1-2z+z=2
-5-z=2
z=-7
substitute x & z in the equation
x + y + 2z = -2
-3+y+2*-7=-2
-3+y-14=-2
y=15

CHECK
x + y + 2z = -2
-3+15+(-14)=-2